OpenStudy (llivingston2):

Multiply (x – 4)(x2 + 5x + 6).

OpenStudy (3mar):

May I help?

OpenStudy (llivingston2):

yes please!!!!!!!

OpenStudy (3mar):

Thanks

OpenStudy (polymorphism):

Take the x and then times it by each term and constant in (x2 + 5x + 6). Also I assume x2 =\(x^2\) So x * x^2 = x^3 x* 5x = 5x^2 x* 6 = 6x So we have x^3 + 5x^2 + 6x Now do the same with -4 -4 * x^2 = -4x^2 -4 * 5x = -20x -4 * 6 = -24 Now put it all together x^3 + 5x^2 + 6x - 4x^2 -20x -24 Which would look like \( x^3 + 5x^2 + 6x - 4x^2 -20x -24\) Can you finish it up from here ?

OpenStudy (llivingston2):

im taking this test right now so im trying to do this fast

OpenStudy (polymorphism):

I am not sure if we are allowed to help with tests

OpenStudy (retireed):

Multiply (x – 4)(x2 + 5x + 6) do it in parts Multiply (x)(x2 + 5x + 6) Multiply ( -4)(x2 + 5x + 6) add the two answers

OpenStudy (llivingston2):

so now i just add all of the like terms

OpenStudy (3mar):